OpenStudy (anonymous):

what is the ratio of a to b 9over 17 = 3a over a+b

OpenStudy (anonymous):

Let's get rid of the denominators

OpenStudy (anonymous):

first we multiply both sides by (a+b), so we get 9/17(a+b) = 3a

OpenStudy (anonymous):

we then multiply both sides by 17, so we get 9(a+b) = 9a+9b=51a

OpenStudy (anonymous):

then subtract 9a from both sides to get 9b=42a

OpenStudy (anonymous):

we find that 42a = 9b, so the ratio of a to b is 42:9

OpenStudy (cggurumanjunath):